# Discrete Autoregressive Switching Processes in Sparse Graphical Modeling of Multivariate Time Series Data

### v1

@@ 4 Application to fMRI Data
Identifying the dynamic nature of brain connectivity is critical for understanding and advancing our current knowledge about human brain functioning. Functional magnetic resonance imaging (fMRI), which measures brain activity by detecting changes associated with blood flow, has become a successful and effective instrument for studying how the brain functions. Here we consider the problem of estimating brain connectivity, i.e., statistical dependence between fMRI time series in distinct regions of the brain. Recent evidence has shown that the interactions among brain regions vary over the course of an fMRI experiment, suggesting that brain connectivity is a dynamic process (Cribben et al.,, 2013; Lindquist et al.,, 2014; Xu and Lindquist,, 2015; Warnick et al.,, 2018). Flexible models that account for dynamic features, change-points and sparse structures are needed for the analysis of such data.
@@
The scanning session proceeded with four blocks consisting of 275 fMRI acquisitions. For the analyses of this paper we concatenated the four blocks and subtracted the mean of each block. $D=18$ lateralized regions of interest (ROIs) were selected on the basis of prior findings using this task (Bornstein and Daw,, 2012, 2013) as those most sensitive to one or more of the identified representations of sequential s tructure (dorsal and ventral striatum, hippocampus), to the degree of conflict between the representations (anterior cingulate cortex), or to the stimulus content (scene images; ventral visual stream regions). We scaled each dimension $d=1,\dots,D$, to have variance one. In our analysis, we seek to identify distinct regimes of functional connectivity that can be mapped onto cognitive interpretation – specifically, to identify the manner in which multiple representations combine to control behavior. Since we do not have prior information on the cognitive states that are manifested during the experiment, we assumed that the number of the states is unknown, as well as the latent learning structure driving the switching of regimes. We set the hyperparameters of the model as described in Section 3.1 and ran MCMC chains with 4000 iterations, 1200 of which were discarded as burn-in.
Figure 5: Application to FMRI data. (top) time series data from one participant, with each dimension represented by a different colored line; vertical colored bands represent the estimated state sequence. (bottom) estimated time-varying state probability plot.
The time series data for the $D=18$ selected ROIs are shown in the top panel of Figure 5, along with the estimated latent state sequence. Our model identified a mode at $\hat{K}=2$ distinct states and a DAR order $\hat{P}=2$, with estimated values of innovations and autoregressive parameters $\hat{\bm{\pi}}=[0.50,0.50]$ and $\hat{\bm{\phi}}=[0.08,0.87,0.05]$, respectively. The bottom panel of Figure 5 displays the time-varying probability plot, namely the local decoding of the hidden state at time $t$, $p(\gamma_{t}=j\,|\,\bm{y},\textperiodcentered)$, for $t=1,\dots,T$, as described in Section 2.4. These probabilities are represented with a different color for each of the two inferred states, and they cumulatively add to one for each $t$. The state probability plot displays a clear transition from state 1 to state 2, approximately halfway through the task.
@@
The top panel of Figure 6 shows the estimated state-specific partial correlation matrices, for the two estimated states, and the bottom panel the corresponding estimated connectivity graphs, with edges identified through the procedure described in Section 2.2. State 1 has relatively stronger connectivity between hippocampus (HC) and anterior cingulate cortex (ACC, ant_cing), with mean difference between states across ROI pairs equal to .028.028.028.028; whereas State 2 shows stronger connectivity between Caudate and ACC, with mean difference between states across ROI pairs equal to .048.048.048.048. Across all ROI pairs, the average difference in partial correlation values between states is .002.002.002.002.
@@ 5 Concluding Remarks
sggmDAR - a Julia software implementing the methodology outlined in the paper, accompanied by a comprehensive tutorial designed to guide users through replicating the findings detailed in the article.
@@ B. Forward Messages for Local Decoding
$\alpha_{t}(j_{1})=\sum_{j_{2}=1}^{\hat{M}}\dots\sum_{j_{\hat{P}}=1}^{\hat{M}}\alpha_{t}(j_{1},j_{2},\dots,j_{\hat{P}}),$ (S.1)
for $j_{l}\in\{1,\dots,\hat{M}\}$, $l=1,\dots\hat{P}$, and the DAR-forward messages $\alpha_{t}(j_{1},\dots,j_{\hat{P}})=p\,(\bm{y}_{1:t-1},\gamma_{t-1}=j_{1},\dots,\gamma_{t-\hat{P}}=j_{\hat{P}}\,|\,\cdot)$ are described as the probability of the partial observations sequence $\bm{y}_{1:t-1}$, and states $\gamma_{t-1:t-\hat{P}}$ at time $t-1$, given all the other parameters. These messages can be recursively computed as follows
@@
Proposition 3. Let $\alpha_{t}(j_{1},\dots,j_{\hat{P}})=p\,(\bm{y}_{1:t-1},\gamma_{t-1}=j_{1},\dots,\gamma_{t-\hat{P}}=j_{\hat{P}}\,|\,\cdot)$ be the DAR-forward messages, described as the probability of the partial observations sequence $\bm{y}_{1:t-1}$, and states $\gamma_{t-1:t-\hat{P}}$ at time $t-1$, given all the other parameters. Then, these messages can be recursively computed as in Eq. (S.2).
@@
C. Simulated scenario with zero-mean observations
@@ D. Large $D$ setting
E. Sensitivity Analysis

### v2

@@ 4 Application to fMRI Data
Identifying the dynamic nature of brain connectivity is critical for understanding and advancing our current knowledge about human brain functioning. Functional magnetic resonance imaging (fMRI), which measures brain activity by detecting changes associated with blood flow, has become a successful and effective instrument for studying how the brain functions. Here we consider the problem of estimating brain connectivity, i.e., statistical dependence between fMRI time series in distinct regions of the brain. Recent evidence has shown that the interactions among brain regions vary over the course of an fMRI experiment, suggesting that brain connectivity is a dynamic process (Cribben et al.,, 2013; Lindquist et al.,, 2014; Xu and Lindquist,, 2015; Vidaurre et al.,, 2017, 2018; Warnick et al.,, 2018). In this context, our proposed modeling framework is particularly well suited to address two key challenges in the analysis of fMRI data. First, the number of distinct cognitive or connectivity states that manifest during the course of the task is unknown and likely to vary across individuals or experimental conditions. Rather than assuming a fixed number of states, our model infers the number of latent regimes in a data-driven fashion using a sparsity-inducing prior. Second, the temporal structure governing transitions between these regimes may involve dependencies that extend beyond first-order Markovian assumptions. Employing a DAR process of unknown order allows for more flexible and biologically plausible switching dynamics. Together, these features enable the identification of meaningful patterns in brain connectivity that evolve over time, offering insight into latent cognitive processes.
@@
The scanning session proceeded with four blocks consisting of 275 fMRI acquisitions. For the analyses of this paper we concatenated the four blocks and subtracted the mean of each block. $D=18$ lateralized regions of interest (ROIs) were selected on the basis of prior findings using this task (Bornstein and Daw,, 2012, 2013) as those most sensitive to one or more of the identified representations of sequential structure (dorsal and ventral striatum, hippocampus), to the degree of conflict between the representations (anterior cingulate cortex), or to the stimulus content (scene images; ventral visual stream regions). The observed time series correspond to the preprocessed and standardized blood-oxygen-level-dependent (BOLD) signals recorded from each of the selected brain regions, serving as proxies for local neural activity across time. We scaled each dimension $d=1,\dots,D$, to have variance one. In our analysis, we seek to identify distinct regimes of functional connectivity that can be mapped onto cognitive interpretation – specifically, to identify the manner in which multiple representations combine to control behavior. Since we do not have prior information on the cognitive states that are manifested during the experiment, we assumed that the number of the states is unknown, as well as the latent learning structure driving the switching of regimes. We set the hyperparameters of the model as described in Section 3.1 and ran MCMC chains with 4000 iterations, 1200 of which were discarded as burn-in.
Figure 6: Application to FMRI data. (top) BOLD time series data from one participant, with each brain region represented by a different line; vertical bands represent the estimated state sequence. (bottom) estimated time-varying state probability plot.
The BOLD time series data for the $D=18$ selected ROIs are shown in the top panel of Figure 6, along with the estimated latent state sequence. Our model identified a mode at $\hat{K}=2$ distinct states and a DAR order $\hat{P}=2$, with estimated values of innovations and autoregressive parameters $\hat{\bm{\pi}}=[0.50,0.50]$ and $\hat{\bm{\phi}}=[0.08,0.87,0.05]$, respectively. The bottom panel of Figure 6 displays the time-varying probability plot, namely the local decoding of the hidden state at time $t$, $p(\gamma_{t}=j\,|\,\bm{y},\textperiodcentered)$, for $t=1,\dots,T$, as described in Section 2.4. These probabilities are represented with a different color for each of the two inferred states, and they cumulatively add to one for each $t$. The state probability plot displays a clear transition from state 1 to state 2, approximately halfway through the task.
@@
The top panel of Figure 7 shows the estimated state-specific partial correlation matrices, for the two estimated states, and the bottom panel the corresponding estimated connectivity graphs, with edges identified through the procedure described in Section 2.2. State 1 has relatively stronger connectivity between hippocampus (HC) and anterior cingulate cortex (ACC), with mean difference between states across ROI pairs equal to .028.028; whereas State 2 shows stronger connectivity between Caudate and ACC, with mean difference between states across ROI pairs equal to .048.048. Across all ROI pairs, the average difference in partial correlation values between states is .002.002. In the Supplementary Material, Section C.2, we provide additional analysis of the fMRI data, including state-specific mean and variance estimates, standardized residual diagnostics, and ACF/PACF plots to further assess model fit and temporal structure. These results show that while mean activation remains largely stable across states, variability differs meaningfully, suggesting distinct dynamic regimes. Residual diagnostics, including standardized residuals and Q-Q plots, further support model adequacy, indicating no major misspecification and approximate normality of the residuals. Together, these findings suggest that our model effectively captures the dominant temporal structure of the data, achieving a balance between flexibility and interpretability in the context of dynamic brain connectivity.
@@ 5 Concluding Remarks
Acknowledgments
B.H.-A. completed part of this work as a postdoctoral fellow in the Department of Statistics at Rice University, Houston, TX.
Disclosure Statement
No potential competing interest was reported by the authors.
@@ Acknowledgments
sggmDAR - a Julia software implementing the methodology outlined in the paper, accompanied by a comprehensive tutorial designed to guide users through replicating the findings detailed in the article. The software sggmDAR is also available on GitHub at https://github.com/Beniamino92/sggmDAR.
SUPPLEMENTARY MATERIAL
This document provides supplemental material for the article “Discrete Autoregressive Switching Processes in Sparse Graphical Modeling of Time Series.” Sections A. and B. present additional details on backward and forward messages, along with the validity of these expressions. Section C. includes supplementary results for the simulation study and fMRI application. Section D. investigates performance under zero-mean emissions, while Section E. examines the approach in high-dimensional settings. Section F. evaluates model performance under misspecification, and Section G. assesses the model’s behavior on synthetic fMRI data. Finally, Sections H. and I. provide sensitivity analysis results and MCMC convergence diagnostics.
@@ B. Forward Messages for Local Decoding
[…]
@@
[…]
@@
C. Additional Results
In this section, we provide supplementary results for the simulation study presented in Section 3 and the application to fMRI data discussed in Section 4.
C..1 Simulation Study under Time-Varying Means
As part of the simulation setting in Section 3, beyond analyzing the estimated state-specific connectivity matrices, it is also informative to examine the state-wise mean and variance values across different dimensions.Figure 1 presents the estimated means (left panel) and variances (right panel) for each state and each variable. The heatmap of mean values reveals distinct patterns across states, with certain variables exhibiting strong shifts in activation levels between states. This suggests that the identified states effectively capture meaningful variations in the underlying data distribution. On the other hand, the variance heatmap indicates that some dimensions exhibit higher variability in specific states, potentially reflecting differences in signal stability or noise contributions across brain regions. Notably, variable $y_{1}$ in State 2 shows a particularly high variance, which may correspond to transient fluctuations in activity or an outlier effect.
Figure 1: Simulation Study. State-specific mean (left) and variance (right) values across different dimensions.
C..2 Application to fMRI Data
As part of the fMRI analysis in Section 4, we report state-specific mean and variance estimates across dimensions in Figure 2. The left panel of the figure shows the estimated means, which remain relatively similar across the two states. In contrast, the right panel illustrates the variance estimates, where State 1 generally exhibits higher variability compared to State 2. This suggests that while the overall activation levels remain stable, the fluctuations in neural activity differ between states.
Figure 2: Application to FMRI data. State-specific mean (left) and variance (right) values across different dimensions.
To investigate the accuracy of the proposed approach in modeling the fMRI data, we display in Figure 3 standardized residuals and corresponding QQ-plots for four representative brain regions. The standardized residuals (left panel) show no obvious pattern, indicating a good fit, while the QQ-plots (right panel) suggest that the residuals closely follow a normal distribution, with only minor deviations in the tails. These results confirm that the model effectively captures the underlying structure of the data.
Figure 3: Application to fMRI Data. (a) Standardized residuals; (b) Q-Q plots of standardized residuals. First four dimensions, representing representative brain regions.
Finally, we display ACF and PACF plots for the same four representative brain regions in Figure 4. These plots provide insight into the temporal structure of the observed fMRI time series, highlighting short-term dependencies that may not always be fully captured by the proposed model. While our approach assumes a discrete autoregressive structure for the evolution of the latent states and independent Gaussian emissions at the observational level, its primary goal is to characterize connectivity patterns through the state-specific precision matrices. Future work could explore extensions that incorporate additional temporal dependencies while preserving the interpretability of the inferred functional connectivity.
Figure 4: Application to fMRI Data. ACF and PACF plots for four representative brain regions, illustrating the temporal dependencies in the observed time series.
D. Simulated scenario with zero-mean observations
@@ E. Large $D$ setting
F. Misspecified Model
We evaluate the performance of our proposed approach under model misspecification. Specifically, we generate data from a stationary vector autoregressive (VAR) process of order 1, defined as
$\bm{y}_{t}=\bm{\mu}+\bm{A}\bm{y}_{t-1}+\bm{\varepsilon}_{t},\qquad\bm{\varepsilon}_{t}\sim N_{D}(\bm{0},\Sigma_{\varepsilon}),$
where $\bm{y}_{t}\in\mathbb{R}^{D}$ represents the $D$-dimensional time series at time $t$, and $\bm{\mu}\in\mathbb{R}^{D}$ is an intercept term that captures the mean level of the process. Here, we generated $T=500$ observations with dimensionality $D=5$. The matrix $\bm{A}\in\mathbb{R}^{D\times D}$ contains the autoregressive coefficients, governing both the temporal dependencies within each variable and the interactions across variables. The innovation term $\bm{\varepsilon}_{t}$ follows a multivariate normal distribution with zero mean and covariance matrix $\Sigma_{\varepsilon}$, capturing the variability and cross-correlations in the residual errors. To ensure stationarity, we assume that the eigenvalues of $\bm{A}$ lie within the unit circle, i.e., all eigenvalues of $\bm{A}$ have modulus strictly less than one. This condition prevents explosive behavior and ensures that past influences decay over time, leading to a stable time series (Lütkepohl,, 2005). We consider two scenarios for the autoregressive coefficient matrix $\bm{A}$: (i) a sparse VAR structure, where only a small subset of entries in $\bm{A}$ are nonzero, reflecting limited and localized interactions among variables. Such sparsity is characteristic of high-dimensional applications like functional MRI (fMRI) data, where neuronal activity exhibits structured but selective dependencies. (ii) A dense VAR structure, where most entries in $\bm{A}$ are nonzero, representing a system with widespread dependencies across variables. This setting corresponds to scenarios where all variables exert direct or indirect influences on each other, such as in economic or climatological models. By analyzing both cases, we assess the robustness of our approach under varying levels of dependency structure, from highly localized interactions to fully connected systems.
Sparse Autoregressive Matrix: The autoregressive coefficient matrix $\bm{A}$ is constructed to be sparse, with a sparsity level of $\rho=0.7$, meaning that only 30% of its entries are nonzero. The nonzero coefficients are drawn uniformly from the interval $0.8,-0.1]\cup[0.1,0.8]$, ensuring a mix of positive and negative dependencies while avoiding near-zero values. The intercept terms $\bm{\mu}$ are sampled from a uniform distribution on [3,3][-3,3], while the covariance matrix $\bm{\Sigma}_{\varepsilon}$ is randomly generated as described in Section 3.1. A realization of the generated time series is shown in Figure 6(d), illustrating the dynamical behavior induced by the sparse structure. Figure 6(a) and Figure 6(e) display the corresponding partial correlation matrix and the sparse autoregressive coefficient matrix, respectively. Notably, the first column of $\bm{A}$ incorporates the intercept $\bm{\mu}$.
We apply our proposed sampler with a maximum allowable number of states set to $M_{\max}=6$, while keeping the remaining parameter settings as described in Section 3. The posterior distribution for the number of states indicates a preference for a single regime, with $P(M=1\mid\cdot)=0.65$ and $P(M=2\mid\cdot)=0.35$. Despite the model misspecification, our approach correctly identifies a dominant regime, yielding an estimated mode at $\hat{M}=1$. Figure 3.1(b) displays the estimated partial correlation matrix, demonstrating strong agreement with the true generating structure. Additionally, Figure 3.1(c) presents the estimated means and variances for each dimension, further supporting the accuracy of our inference. These results highlight the robustness of our method in capturing the underlying dependency structure, even when the data-generating process deviates from the assumed model. Notably, in the case of a sparse autoregressive matrix, the method predominantly identifies a single latent state, suggesting that sparsity in the autoregressive structure naturally reduces regime complexity.
Figure 6: Misspecified Model - Sparse VAR. (a) true partial correlation matrix; (b) estimated partial correlation matrix; (c) estimated mean and variances; (d) time series realization; (e) generated VAR matrix.
Dense Autoregressive Matrix: We now consider the case of a dense autoregressive coefficient matrix $\bm{A}$, where the sparsity level is set to $\rho=0.1$, meaning that 90% of its entries are nonzero. The nonzero coefficients and the intercept term $\bm{\mu}$ are generated following the same procedure as in the sparse case. A realization of the generated time series is shown in Figure 7 (top), illustrating the more intricate dependencies introduced by the dense structure. Figures 8(b) and 8(c) display the corresponding partial correlation matrix and dense autoregressive coefficient matrix, respectively. The remaining parameterization for the proposed model is kept identical to the sparse case.
In this setting, the posterior distribution for the number of states indicates a strong preference for three regimes, with $P(M=2\mid\cdot)=0.01$ and $P(M=3\mid\cdot)=0.99$, yielding an estimated mode at $\hat{M}=3$. Thus, under model misspecification with a dense autoregressive structure, our approach identifies multiple regimes, in contrast to the single-state solution found in the sparse case. Figure 9 presents the time-varying probability plot, illustrating the inferred state transitions over time. Additionally, Figure 8(a) shows the estimated state-specific partial correlation matrices, which align closely with the true generating matrix, confirming that the model captures key structural dependencies.
To further examine the factors driving the differentiation into three distinct states, we analyze the state-specific means and variances across dimensions, as shown in Figure 8(d). Notably, the variance in state 2 for components 3 and 4 differs markedly from the other states, while component 4 also exhibits noticeable shifts in its mean. These distinctions suggest that, in the presence of a dense autoregressive structure, variability in both the mean and variance plays a critical role in defining latent regimes.
Finally, to assess the model’s ability to capture the autocorrelation structure of the data, despite the misspecification, we examine the standardized residuals and QQ plots in Figure 9. The results indicate an overall satisfactory goodness of fit, providing further evidence that our approach effectively extracts meaningful temporal patterns even when the assumed model deviates from the true data-generating process.
Figure 7: Misspecified Model - Dense VAR. (top) time-series realization (lines) where each dimension is represented by a different colored line; (bottom) estimated time-varying probability plot.
Figure 8: Misspecified Model - Dense VAR (a) estimated state-specific partial correlation matrices; (b) true partial correlation matrix; (c) generated VAR matrix; (d) estimated mean and variances.
Figure 9: Misspecified Model - Dense VAR. (a) standardized residuals; (b) Q-Q plots of standardized residuals.
G. Simulated fMRI data
In this section, we investigate the performance of our proposed model by evaluating its behavior under a controlled simulation framework. We generate synthetic fMRI data using the neuRosim package in R, which provides a principled approach to simulating BOLD signals, stimulus-driven activation, and multiple sources of structured noise. The simulation mimics a typical block-design fMRI experiment with region-specific activation patterns and realistic noise components.
G..1 BOLD Signal Generation
The synthetic dataset consists of $D=10$ brain regions and $T=200$ time points, corresponding to a total scan duration of 400 seconds with a repetition time (TR) of 2 seconds. The experimental design follows a block paradigm where an external stimulus is periodically introduced every 80 seconds, lasting for 40 seconds. The stimulus function is generated using the stimfunction() function from neuRosim, creating a high-resolution representation of the experimental design. This function is subsequently downsampled to match the fMRI temporal resolution. For each brain region, the true underlying BOLD signal is simulated using specifydesign(), which convolves the stimulus function with a canonical double-gamma hemodynamic response function (HRF). The effect size of activation varies across regions, ensuring heterogeneous responses to the stimulus. Figure 10 (middle) displays the generated true underlying BOLD signal.
To model real-world fMRI data, we introduce multiple noise sources, each uniquely affecting individual brain regions. Specifically, (i) temporal noise is modeled as an auto-regressive (AR) process with parameters $\rho_{1}=0.5,\rho_{2}=-0.3$, generated using temporalnoise(), accounting for slow signal drifts and temporal autocorrelation typical of fMRI data; (ii) physiological noise, induced by cardiac and respiratory cycles, is introduced using physnoise(), which generates region-specific fluctuations that exhibit structured oscillatory patterns influenced by the fMRI repetition time; and (iii) scanner noise is incorporated via systemnoise(), modeling Rician-distributed noise to account for thermal and hardware-induced variability in the fMRI acquisition process. Each region is assigned a noise level sampled from a uniform distribution between 1 and 3, reflecting spatially heterogeneous scanner noise. The final data representation is obtained by summing the clean BOLD signal with the noise components and standardizing each brain region by subtracting its mean and dividing by its standard deviation, as shown in Figure 10 (top). By introducing region-specific variability across all noise sources, the simulated dataset closely resembles real fMRI time series, providing a robust testbed for evaluating the proposed methodology.
G..2 Results
We apply our proposed sampler with the maximum allowable number of states set to $M_{\max}=6$, while keeping the remaining parameter settings consistent with those described in Section 3. The posterior distribution on the number of states exhibits a strong mode at $\hat{M}=2$, with an overwhelming posterior probability of $p(M=1\mid\cdot)=0.99$, indicating that the model strongly favors a two-state structure. Similarly, the posterior distribution on the order of the discrete autoregressive (DAR) process was estimated at $\hat{P}=2$, with posterior probabilities $p(P=1\mid\cdot)=0.15$, $p(P=2\mid\cdot)=0.79$, and $p(P=3\mid\cdot)=0.06$, confirming the adequacy of modeling dependencies up to two previous time points. Figure 10 (bottom) displays the time-varying probability plots inferred by the model. These plots reveal a clear alignment between the estimated state transitions and the periodic stimulus changes embedded in the true generated BOLD signal. This suggests that the proposed methodology effectively captures the latent switching dynamics in an unsupervised manner.
To further assess the characteristics of the inferred states, we examine the estimated state-specific means and variances, as shown in Figure 11. The heatmaps illustrate distinct mean activation patterns across the two states, with one state exhibiting consistently higher activation levels across most brain regions, while the other reflects lower or baseline activation. This separation of states suggests that the model successfully distinguishes between task-evoked and resting-state activity. The variance estimates, presented in the right panel of Figure 11, indicate that variability remains relatively stable across most regions, though with some fluctuations in specific regions, notably in region $y_{7}$. This might reflect localized differences in the neural response to the stimulus or variations in noise contributions across regions.
Finally, we evaluate the model’s goodness-of-fit by analyzing standardized residuals. Figure 12(a) presents the residual time series for four selected brain regions, while the corresponding QQ plots in Figure 12(b) assess their adherence to a normal distribution. While the residuals generally align well with normality, some deviations in the tails suggest minor departures, which may indicate room for further refinement in capturing noise structure. Nevertheless, the overall results confirm that the model successfully detects state transitions, accurately characterizes the stimulus-driven dynamics, and provides meaningful statistical summaries of the two inferred states.
Figure 10: Simulated fMRI Data. (a) Simulated fMRI data obtained by summing the clean BOLD signal with noise components and standardizing each brain region; (b) True underlying BOLD signal, generated by convolving the stimulus function with a canonical double-gamma HRF; (c) Estimated time-varying probability plot.
Figure 11: Simulated fMRI data. State-specific mean (left) and variance (right) values across different dimensions.
Figure 12: Simulated fMRI data. (a) standardized residuals; (b) Q-Q plots of standardized residuals.
H. Sensitivity Analysis